Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Wavefront sensing)
lumaca, nardecchia, rocchi, berni, pinto, ruggi, spinicelli - 23:05 Monday 29 May 2023 (60362) Print this report
WI absorption measurement-part 2

This afternoon, other two ‘short’ measurements have been performed to observe the PAs:

1) In gentle unlock configuration:
     -14.29 UTC: HWS INJ acquisition started (20230529T1629)
    -14.37 UTC:  Gentle Unlock (not worked) from low noise.
                         CH flip mirror manually closed  (WF 15).
                          Francesco realigned PR and WI at 14.38 UTC (WF 17).

      -15.20 UTC: HWS measurement interrupted (WF 95)
      -15.21 UTC: CH flip mirrors opened.

Info from Manuel and Francesco: the gentle unlock can be perform only starting from CARM NULL 1F

      2) During the lock acquisition:
            - 17.04 UTC: HWS INJ acquisition started IN CITF (20230529T1904)
            - 17.10 UTC: start lock acquisition  WF 11
            -17.12 UTC:  CH switched off   WF 14             
            - 17.12 UTC: CARM NULL 3F WF 15
           - 17.14 UTC: CARM NULL 1F WF 18
           - 17.22.59 UTC : Unlock during the OMC locking phase (WF33)
The data will be carefully analyzed to check the effect of the changing alignment.

Comments to this report:
nardecchia - 19:27 Tuesday 11 July 2023 (60798) Print this report

Concerning the measurement performed during the lock acquisition, I re-run the analysis by using different wavefronts as reference.

Here the lists of all steps in detail:

  1. the lock acquisition started at 17.10.30 UTC

  2. first wavefront after starting the lock acquisition: WF 11 (17.12.08 UTC)

  3. CH OFF at  17.11.57 UTC, WF 14 acquired at 17.12.08 UTC

  4. WF 15 acquired at 17.12.44 UTC

  5. CARM NULL 3F acquired at  17.12.48 UTC 

  6. WF 16 acquired at  17.13.14 UTC

  7. ITF unlock at 17.22.56 (WF 33)

The data have been re-processed considering the wavefronts in the range between 11 and 17 but, only the differential wavefronts computed wrt the images #17 look reliable.

An example is shown in fig.1 where both YAG and point absorbers OPLs are shown.

The WF17 has been acquired when the ITF lock index was 104.

Note that by re-analyzing the data considering the previous wavefront , WF16, acquired at ITF lock index 100, you can not see anything.

Madda told me that 100 corresponds to CARM NULL 3F and 104 to CARM null 1F

Thus, to perform HWS measurements during the lock acquisition, it is needed to take the reference at CARM null 1f (ITF LOCK INDEX 104).

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×